OpenStudy (anonymous):

W X Y Z is a number bar if W+Z=X+Y and 2W=3X If W X Y W is a number bar, Y=

OpenStudy (anonymous):

any tips?

OpenStudy (aripotta):

since we know that 2w = 3x, we put 3x/2 into the equation to substitute for w so, 3x/2+z=x+y then, we can multiply 2 on each side 2[(3x/2)+z]=2(x+y) 3x+2z=2x+2y then subtract 2x from each side x+2z=2y divide both sides by 2 y=x+2z/2

OpenStudy (anonymous):

yeah

OpenStudy (aripotta):

is that one of the answers? lol

OpenStudy (anonymous):

but what would W equal?

OpenStudy (anonymous):

3/4W, W, 4/3W, 3W, or 4W?

OpenStudy (aripotta):

oh i thought you were just finding y

OpenStudy (anonymous):

Does Y equal 3/4W, W, 4/3W, 3W, or 4W?

OpenStudy (aripotta):

O-o lemme think

OpenStudy (anonymous):

ty for helping me; I really appreciate it.

OpenStudy (anonymous):

2(2/3W=4/3W?

ganeshie8 (ganeshie8):

y = (4/3)W

OpenStudy (aripotta):

how'd you get that?

ganeshie8 (ganeshie8):

W X Y W 2W = X + Y 2W = (2/3)W + Y solve for Y

OpenStudy (anonymous):

Let us do a Geometry problem now

ganeshie8 (ganeshie8):

yea

OpenStudy (aripotta):

derp.

ganeshie8 (ganeshie8):

DERP means "Reply to stupid comment or action"

OpenStudy (aripotta):

my stupid reply lol. i'm derping myself

ganeshie8 (ganeshie8):

^ lol